OpenStudy (gavin39):

.What polynomial has roots of -5, -4, and 1 ? x3 - 8x2 - 11x + 20 x3 - x2 + 22x + 40 x3 + x2 - 22x - 40 x3 + 8x2 + 11x - 20

OpenStudy (gavin39):

@StudyGurl14

OpenStudy (gavin39):

Will give medals!

OpenStudy (gavin39):

@Mertsj

OpenStudy (studygurl14):

I forgot how to do this...

Directrix (directrix):

If -5 is a root, then (x -(-5)) = (x + 5) is a factor. If -4 is a root, then (x - (-4)) = (x + 4) is a factor. If 1 is a root, the (x - 1) is a factor.

Directrix (directrix):

To get the polynomial, multiply these 3 factors: (x + 5)*(x +4)*(x - 1) = ? @gavin39